OpenStudy (anonymous):

Medal, Fan, and Testimonial! suppose that y varies inversely with x, and y = 4 when x=12. What is the equation for the inverse variation?

OpenStudy (anonymous):

This is for a study guide, not for a test or lesson or anything like that. I need to know how to do it, not just the answer. @BloomLocke367 @iGreen

OpenStudy (anonymous):

I have a few more also. Please help! :)

OpenStudy (damonte212):

Inverse Variation Formula: y = k / x Substitute for x and for y: = k / Solve for k: = k New equation: y = / x Substitute for x: y = / Answer: y =

OpenStudy (damonte212):

can you get it that way? or do you want the answer?

OpenStudy (anonymous):

Okay, so it would be 4 = k/12

OpenStudy (damonte212):

correct

OpenStudy (anonymous):

Then it would be k = 48

OpenStudy (damonte212):

right

OpenStudy (anonymous):

48/12 would be 4. So 4?

OpenStudy (anonymous):

Thanks. I thought I knew how to do it, but couldn't remember the formula! 0.0

OpenStudy (damonte212):

np